OpenStudy (anonymous):

What is the simplified form of:

OpenStudy (anonymous):

\[(\frac{ 24y^5 }{ 15x^8 })\div(\frac{ 8y^2 }{ 4x^4 })\]

OpenStudy (anonymous):

@nukka

OpenStudy (anonymous):

((24y^5)/(15x^8))/((8y^2)/(4x^4)) = ((24y^5)/(15x^8))*((4x^4)/(8y^2)) =(24*4)(x^4)(y^5)/(15*8)(x^8)(y^2) =(4/5)(1/x^4)(y^3)